# Kolmogorov-Arnold Energy Models: Fast and Interpretable Generative Modeling

### v8

@@ A.4.7 Learning gradient for discretized Thermodynamic Integral and power posterior partition function
A.5 Sampling theory
This section provides more theoretical detail regarding KAEM’s sampling procedures.
A.5.1 Importance Sampling
Importance Sampling (IS) is a Monte Carlo that enables the estimation of expectations with respect to one probability measure using samples drawn from another. For a detailed exposition, see 910e8e7d-2e70-317c-adf6-e21cd50755a1.
The Radon-Nikodym derivative
IS involves rewriting the expectation of an arbitrary function or density, $\rho(\bm{z})$, under the target measure, $\mathbb{P}(\bm{z})$, in terms of a more tractable proposal measure, $q(\bm{z})$. Specifically, we use the Radon-Nikodym derivative $\frac{d\mathbb{P}}{d\mathbb{Q}}$ to rewrite the expectation of an arbitrary function of $\bm{z}$, denoted as $\rho(\bm{z})$:
$\mathbb{E}_{\mathbb{P}(\bm{z})}[\rho(\bm{z})]=\int_{\mathcal{\bar{Z}}}\rho(\bm{z})\,\mathbb{P}(d\bm{z})=\int_{\mathcal{\bar{Z}}}\rho(\bm{z})\cdot\frac{d\mathbb{P}}{d\mathbb{Q}}\,\mathbb{Q}(d\bm{z}).$ (59)
In the context of KAEM, the target distribution is the posterior, $p_{\bm{f},\bm{\Phi}}(\bm{z}\mid\bm{x}^{(b)})$, which is proportional to the product of the likelihood and the prior:
$p_{\bm{f},\bm{\Phi}}(\bm{z}\mid\bm{x}^{(b)})\propto p_{\bm{\Phi}}(\bm{x}^{(b)}\mid\bm{z})\cdot p_{\bm{f}}(\bm{z}).$ (60)
The exact normalization constant, (the marginal likelihood $p_{\bm{f},\bm{\Phi}}(\bm{x}^{(b)})$), is intractable, but Importance Sampling circumvents the need to compute it explicitly. Instead, the expectation of $\rho(\bm{z})$ under the posterior can be expressed as:
$\mathbb{E}_{p_{\bm{f},\bm{\Phi}}(\bm{z}\mid\bm{x}^{(b)})}[\rho(\bm{z})]=\int_{\mathcal{\bar{Z}}}\rho(\bm{z})w(\bm{z})\,\mathbb{Q}(d\bm{z}),$ (61)
where the importance weights $w(\bm{z})$ are given by:
$w(\bm{z})=\frac{p_{\bm{f},\bm{\Phi}}(\bm{z}\mid\bm{x}^{(b)})}{q(\bm{z})}=\frac{p_{\bm{\Phi}}(\bm{x}^{(b)}\mid\bm{z})\cdot p_{\bm{f}}(\bm{z})}{q(\bm{z})}.$ (62)
Proposal distribution and importance weights
A practical choice for the proposal density $q(\bm{z})$ is the prior, $p_{\bm{f}}(\bm{z})$, given the availability of draws from the prior using the method outlined in Sec. 2.4.3. This simplifies the importance weights, as the prior has a tractable form and covers the support of the posterior. This yields:
$w(\bm{z}^{(s)})=p_{\bm{\Phi}}(\bm{x}^{(b)}\mid\bm{z}^{(s)})\propto\exp(\log(p_{\bm{\Phi}}(\bm{x}^{(b)}\mid\bm{z}^{(s)}))).$ (63)
This reflects that the likelihood now directly informs how much weight to assign to each sample $\bm{z}$. Intuitively, samples that better explain the observed data $\bm{x}^{(b)}$, (as measured by the likelihood), are given higher importance. Using the definition of softmax, the normalized weights are given by:
$\frac{w(\bm{z}^{(s)})}{\sum_{r=1}^{N_{s}}w(\bm{z}^{(r)})}=\text{softmax}_{s}\Big(\log(p_{\bm{\Phi}}(\bm{x}^{(b)}\mid\bm{z}^{(s)}))\Big).$ (64)
Resampling
The suitability of these weights depends on how well $q(\bm{z})$ matches the posterior. In KAEM, there is likely to be significant mismatch due to the complexity of the likelihood models described in Eq. 20. Consequently, the weights will exhibit high variance, resulting in many samples contributing minimally. This reduces the Effective Sample Size (ESS) and introduces bias into the estimates.
To address this, the weights are resampled before proceeding with the posterior expectation. This is accomplished by using one of the methods outlined by douc2005comparisonresamplingschemesparticle, whenever the ESS falls below a certain threshold determined by Eq. 26. Resampling redistributes the sample weights, creating a new population with uniformly distributed weights while preserving the statistical properties of the original distribution. As training progresses and the prior becomes a better match for the posterior, the need for resampling decreases.
Importance Sampling estimator
To estimate the expectation, we draw $N_{s}$ independent samples $\{\bm{z}^{(s)}\}_{s=1}^{N_{s}}\sim q(\bm{z})$ using the procedure outlined in Sec. 2.4.3 and compute a weighted sum:
$\mathbb{E}_{p_{\bm{f},\bm{\Phi}}(\bm{z}\mid\bm{x}^{(b)})}[\rho(\bm{z})]$ $\approx\sum_{s=1}^{N_{s}}\rho(\bm{z}^{(s)})\cdot w_{\text{norm}}(\bm{\bm{z}^{(s)})}$ (65)
$\approx\frac{1}{N_{s}}\sum_{s=1}^{N_{s}}\rho(\bm{z}^{(s)}_{\text{resampled}})$. (66)
The estimator is unbiased, implying that its expectation matches the true posterior expectation. As $N_{s}\to\infty$, the approximation converges almost surely to the true value, provided that $q(\bm{z})$ sufficiently covers the posterior to ensure a well-conditioned Radon-Nikodym derivative. This estimator is used for the log-marginal likelihood and its gradient in Eqs. 15 &  16, since it provides a means of estimating expectations with respect to the posterior distribution.
A.5.2 Residual resampling
In this study, we adopted residual resampling to filter the population of latent samples whenever the effective sample size (ESS) criterion in Eq. 26 is satisfied. Given a set of normalized weights $\{w_{\text{norm}}(\bm{z}^{(s)})\}_{s=1}^{N}$, the method deterministically replicates samples according to the integer part of $Nw_{\text{norm}}$, and uses multinomial sampling to allocate the remaining draws according to the residual mass.
  1. 1.
    Integer replication: Each sample $s$ is deterministically replicated
    $r_{s}=\left\lfloor N\,w_{\text{norm}}(\bm{z}^{(s)})\right\rfloor,$ (67)
    resulting in $\sum_{s=1}^{N}r_{s}$ allocated indices.
  2. 2.
    Residual weights: The number of remaining draws is
    $R=N-\sum_{s=1}^{N}r_{s}.$ (68)
    The unallocated probability mass (fractional remainder) is computed as
    $m_{s}=N\,w_{\text{norm}}(\bm{z}^{(s)})-r_{s}.$ (69)
    If $R>0$, the residual weights are obtained by normalizing this mass:
    $w_{\text{residual}}(\bm{z}^{(s)})=\frac{m_{s}}{R},$ (70)
    otherwise $w_{\text{residual}}(\bm{z}^{(s)})=0$ for all $s$.
  3. 3.
    Multinomial allocation: The remaining $R$ indices are sampled according to the cumulative distribution function of the residual weights. For each draw,
    $k^{*}=\min\left\{j\mid\sum^{j}_{s=1}w_{\text{residual}}(\bm{z}^{(s)})\geq u_{i}\right\}\quad\text{where}\quad u_{i}\sim\mathcal{U}([0,1]).$ (71)
    The final resampled index set consists of the deterministic replications given by $\{r_{s}\}$ and the stochastic residual allocations drawn from Eq. 71. After resampling, weights are reset to a uniform distribution and a standard Monte Carlo estimator can be used.
A.6 Hyperparameters and training details
Table 4: Hyperparameters for Sec. 3.1.1.
@@ A.6 Hyperparameters and training details
Table 5: Hyperparameters for Sec. 3.2.
@@
layer_widths $p,q$ See Sec. A.6.1
spline_type (-) RBF
spline_activation (-) Relu
@@
grid_size $N_{\text{grid}}$ 20
@@
Initial grid_range (before grid updating) $z_{\text{min}},z_{\text{max}}] [$1,1][-1,1]
@@
generator_noise $\sigma_{\epsilon}$ 0.1
generator_variance $\sigma_{\epsilon}$ 0.3
@@
update_prior_grid (-) true
@@
num_generated_samples (-) 3,000
Number of epochs $N_{\text{epochs}}$ 40
@@
A.6.1 SVHN & CelebA Architectures
Table 6: The dense RBF KAN network used as the mixture prior model

### v9

@@ E.0.7 Learning gradient for discretized Thermodynamic Integral and power posterior partition function
Appendix F Supplementary experiment details
F.1 Hyperparameters tables
Table 4: Hyperparameters for NIST datasets in Sec. 4.2.
@@ F.1 Hyperparameters tables
Table 5: Hyperparameters for SVHN and CelebA in Sec. 4.2.
@@
layer_widths $p,q$ See Sec. 3.1
spline_type (-) Morlet Wavelet
spline_activation (-) None
@@
grid_size $N_{\text{grid}}$ N/A
@@
Initial grid_range (before grid updating) $z_{\text{min}},z_{\text{max}}] [$1.5,1.5][-1.5,1.5]
@@
generator_noise $\sigma_{\epsilon}$ 0.25
generator_variance $\sigma_{\epsilon}$ 0.25
@@
update_prior_grid (-) false
@@
num_generated_samples (-) 20,000
Number of epochs $N_{\text{epochs}}$ 200
@@
F.2 SVHN & CelebA Architectures
F.2.1 KAEM Architectures
Table 6: The dense Morlet wavelet KAN network used as the mixture prior model
@@ 5 $\times4$ convT(3), Sigmoid $64\times 64\times 3$ 2
F.2.2 Baseline VAE architectures
Table 9: VAE Encoder Architecture for SVHN ($32\times 32$)
Layer Operation Output Shape Stride
Input Image $\bm{x}$ $32\times 32\times 3$
1 $\times4$ conv($64$), LeakyReLU $16\times 16\times 64$ 2
2 $\times4$ conv($128$), LeakyReLU $8\times 8\times 128$ 2
3 $\times4$ conv($256$), LeakyReLU $4\times 4\times 256$ 2
4 Flatten + FC $\bm{\mu},\log\bm{\sigma}^{2}\in\mathbb{R}^{81}$
Config: kernels=4,$4,$ strides=2/2/2, paddings=1/1/1, latent_dim=81, batchnorm=False
Table 10: VAE Decoder Architecture for SVHN ($32\times 32$)
Layer Operation Output Shape Stride
Input Latent vector $\bm{z}$ $81$
1 FC + Reshape $4\times 4\times 256$
2 $\times4$ convT($128$), LeakyReLU $8\times 8\times 128$ 2
3 $\times4$ convT($64$), LeakyReLU $16\times 16\times 64$ 2
4 $\times4$ convT(33), Sigmoid $32\times 32\times 3$ 2
Config: kernels=4,$4,$ strides=2/2/2, paddings=1/1/1, latent_dim=81, batchnorm=False
Table 11: VAE Encoder Architecture for CelebA ($64\times 64$)
Layer Operation Output Shape Stride
Input Image $\bm{x}$ $64\times 64\times 3$
1 $\times4$ conv($64$), LeakyReLU $32\times 32\times 64$ 2
2 $\times4$ conv($128$), LeakyReLU $16\times 16\times 128$ 2
3 $\times4$ conv($256$), LeakyReLU $8\times 8\times 256$ 2
4 $\times4$ conv($512$), LeakyReLU $4\times 4\times 512$ 2
5 Flatten + FC $\bm{\mu},\log\bm{\sigma}^{2}\in\mathbb{R}^{81}$
Config: kernels=$4,4,$ strides=2/2/2/2, paddings=1/1/1/1, latent_dim=81, batchnorm=False
Table 12: VAE Decoder Architecture for CelebA ($64\times 64$)
Layer Operation Output Shape Stride
Input Latent vector $\bm{z}$ $81$
1 FC + Reshape $4\times 4\times 512$
2 $\times4$ convT($256$), LeakyReLU $8\times 8\times 256$ 2
3 $\times4$ convT($128$), LeakyReLU $16\times 16\times 128$ 2
4 $\times4$ convT($64$), LeakyReLU $32\times 32\times 64$ 2
5 $\times4$ convT(33), Sigmoid $64\times 64\times 3$ 2
Config: kernels=$4,4,$ strides=2/2/2/2, paddings=1/1/1/1, latent_dim=81, batchnorm=False
Appendix G Generated images
G.1 MNIST & FMNIST
(a) Real samples
(b) No ref. prior
(c) Gaussian ref. prior
(d) Uniform ref. prior
Figure 13: Generated MNIST samples
(a) Real samples
(b) No ref. prior
(c) Gaussian ref. prior
(d) Uniform ref. prior
Figure 14: Generated FMNIST samples
G.2 SVHN & CelebA
(a) Real samples
(b) VAE baseline
(c) MLE-trained with ULA sampling
(d) Thermo-trained with population-based ULA
Figure 15: Generated SVHN samples
(a) Real samples
(b) VAE baseline
(c) MLE-trained with ULA sampling
(d) Thermo-trained with population-based ULA
Figure 16: Generated CelebA samples
Appendix H Smaller RBF Experiment
To investigate how hyperparameter choice affects sample quality, we conducted a smaller-scale supplementary experiment using RBF bases. Most hyperparameters were kept consistent with Tab. 5, except for using RBF bases with ReLU base activation and 20 grid points, alongside a smaller additive noise variance of $\sigma_{\text{noise}}=0.1$ and a larger posterior sampling step size, $\eta=0.01$. Training was conducted for 40 epochs with 20,000 training examples and a batch size of $N=50$ (i.e., only 16,000 parameter updates).
The grid updating procedure proposed by Liu et al. (2024) was applied periodically, starting at once every 100 parameter updates with a frequency decay of 0.999, using 100 samples to update the prior’s domain based on regions explored by ULA.
Effectively unbiased metrics were collected using regression on sample sizes 1000, 1200, 1400, 1600, 1800, 2000, 2200, 2400, 2600, 2800, 3000 (i.e., significantly smaller than those used in the main experiment). Regression statistics, (standard errors and $R^{2}$), were not collected for this supplementary experiment.
Model $\overline{\text{FID}}_{\infty}$ $\overline{\text{KID}}_{\infty}$
MLE 78.41 0.0627
Thermo 165.79 0.1608
Table 13: SVHN ($32\times 32$), RBF experiment.
Model $\overline{\text{FID}}_{\infty}$ $\overline{\text{KID}}_{\infty}$
MLE 130.05 0.1427
Thermo 150.25 0.1637
Table 14: CelebA ($64\times 64$), RBF experiment.
(a) Real samples
(b) Maximum Likelihood
(c) Steppingstone Estimator
Figure 17: Generated SVHN ($32\times 32$) samples using RBF bases with mixture prior architecture
(a) Real samples
(b) Maximum Likelihood
(c) Steppingstone Estimator
Figure 18: Generated CelebA ($64\times 64$) samples using RBF bases with mixture prior architecture